Given f(x) = ax3 + 2x2 + bx − 20 where x − 2 and x − 5 are two of its factors determine the real values of a and b


1
Expert's answer
2021-11-07T16:57:54-0500

Given,

"f(x) = ax^3 + 2x^2 + bx \u2212 20"

Factor x − 2 and x − 5

Now, substituting the value of x=2 and x =5 in the above equation

"ax^3+2x^2+bx-20=0"

x=2

"8a+8+2b-20=0"

Now,

"8a+2b=12"

"\Rightarrow 4a+b=6...(i)"

Similarly,

x=5

"125a+50+5b-20=0"

"\Rightarrow 125a+5b+30=0"

"\Rightarrow 25a+b=-6...(ii)"

From equation (i) and (ii)

"21a=-12"

"a=\frac{-4}{7}"

Now, substituting the value a in the above equation,

"b=6-\frac{16}{7}"


"\Rightarrow b= \frac{42-16}{7}"

"\Rightarrow b = \frac{26}{7}"